Created: 07 February 2020
As residents will be aware, Andrew is continuing to monitor and raise the on-going unacceptable issues with the Roxby Landfill.
Andrew explains, "I have received a further update from the Environment Agency (EA) today, which I have copied below for your convenience. I will also be driving to Winterton and Roxby today to once again inspect the odours directly. Additionally, I will be meeting with the EA next Friday for further discussions on this issue. As always, I will be sure to keep residents informed as I receive further information."
Andrew says, "This afternoon, I have also requested that the EA formally suspend the permit and halt all further waste until these problems are resolved. With the further breaches being identified and the fact that residents have been physically sick, this is now beyond a joke. I am not sure if the EA have the power to do so at this stage; however, as mentioned above, I will be meeting the EA next week to discuss further."
Environment Agency Update
Between 3 December 2019 and 6 February 2020 the EA have received 297 odour reports. The attached graph shows the odour reports since 3 December 2019.
Most recently an Environment Agency team was deployed on 6 February 2020 to assess the latest peak in reports of odour. Further breaches of the environmental permit were identified. Biffa had issues with cracking in the temporary capping on the south east corner of cell 9/11. This has been caused by the construction of a temporary haul road in the vicinity. The EA consider these cracks to be an additional source of odour and are contributing to the pollution detected off site. Further works have been identified to address these specific issues.
This new work is in addition to the work that Biffa committed to as part of their voluntary action plan to reduce off-site odours. Activities against the action plan are proceeding in line with the agreed timescale. Fifteen deep gas wells have been installed and are under extraction. Pin Wells are now in place across the site’s northern flank and a large area (approximately 7700m2) of site has been temporarily capped.
The EA continue to monitor the situation closely; undertaking regular site inspections and off-site surveys for assessment of odour in the locality. The EA want to reassure residents that actions are being taken by the operator to resolve the issues and therefore minimise the impact the operations are having on the local environment.